# Improving Human-AI Coordination through Adversarial Training and Generative Models

### v2

@@ 7 Conclusion
Our work demonstrates that combining generative modeling with adversarial training provides an effective approach for training AI agents that can coordinate with diverse human partners. By using a generative model to constrain the Adversary to valid cooperative behaviors while employing regret-based optimization to create increasingly challenging scenarios, we enable robust learning without the risk of degenerate strategies. Experimental results on the Overcooked benchmark show that GOAT achieves state-of-the-art performance in both simulated evaluations and studies with real human partners. The improved coverage of the partner space demonstrated by our agents suggests this approach successfully bridges the gap between training efficiency and the generalization to human behavior. Future work could explore extending this framework to more complex collaborative tasks and investigating ways to incorporate explicit human feedback during training. From a broader perspective, this work represents a step toward AI systems that can flexibly coordinate with humans in real-world settings.
Broader Impact Statement
@@ 8 Acknowledgements
This work was generously supported by the Amazon+UW Science Hub, the Cooperative AI Foundation, and the Sony Research Award Program.
Supplementary Materials
The following content was not necessarily subject to peer review.
9 Reproducibility
We have our demo, code, and extended experimental data on this webpage. https://sites.google.com/view/human-ai-goat-2025. We provide all the details to setup the environment and reproduce the results.
10 Implementation details
Generative models. The VAE model was trained on a dataset containing joint trajectories of two players. This dataset was created by evenly sampling from the pairs in the simulated agent population $\{\pi_{1},...,\pi_{N}\}$ generating 100k joint trajectories. The dataset was split in a 70/30 format for training and evaluation. The VAE was trained using ELBO loss, and linear scheduling of the KL penalty coefficient $\beta$ was applied to control the KL divergence of the posterior distribution. Automatically, checkpoints of VAE models of different strengths were saved.
Cooperator Agent The Cooperator agent was trained using PPO (Schulman et al., 2017). To promote exploration, the first 100M steps use reward shaping for dish and soup pick-up. We adapted our codebase on the previous implementation of HSP (Yu et al., 2022) and GAMMA (Liang et al., 2024).
GOAT For the GOAT policy training, we usethe REINFORCE objective with mean and standard deviation bounds to keep the policy constrained in the VAE latent space. This objective requires two separate passes of the environment to gather returns for self-play and cross-play.
11 Hyperparameters
@@
The architecture of the policy model and the generative model are comparable. The representations are transformed into variational posterior and action reconstruction predictions using an encoder head and a decoder head.

### v3

@@ 7 Conclusion
Our work demonstrates that combining generative modeling with adversarial training provides an effective approach for training AI agents that can coordinate with diverse human partners. Using a generative model and regret-based optimization to constrain the Adversary enables GOAT to generate valid, meaningful cooperation partners while posing a curriculum of increasingly challenging scenarios, adapted to the weaknesses of the Cooperator’s policy. Thus, we enable robust learning without the risk of degenerate strategies. Experimental results on the Overcooked benchmark show that GOAT achieves state-of-the-art performance in both simulated evaluations and studies with real human partners. The improved coverage of the partner space demonstrated by our agents suggests this approach successfully bridges the gap between training efficiency and the generalization to human behavior. Future work could explore extending this framework to more complex collaborative tasks and investigating ways to incorporate explicit human feedback during training.
@@ 8 Acknowledgements
This work was generously supported by the Cooperative AI Foundation, UW-Amazon Science Gift Hub, Amazon Middle Mile Products and Technology Team, Sony Research Award, UW-Tsukuba Amazon NVIDIA Cross Pacific AI Initiative (XPAI), and Microsoft Accelerate Foundation Models Research Program.
Appendix A Reproducibility
We have our live interactive demo and extended experimental data on this webpage. https://sites.google.com/view/goat-2025. We provide all the details to set up the environment and reproduce the results. For the One-Step Cooperative Matrix Game please find Google Colab with code and all the instructions to run the code.
Appendix B Implementation details
Generative models. The VAE model was trained on a dataset containing joint trajectories of two players. This dataset was created by evenly sampling from the pairs in the simulated agent population $\{\pi_{1},...,\pi_{N}\}$, generating 100k joint trajectories. The dataset was split in a 70/30 format for training and evaluation. The VAE was trained using ELBO loss, and linear scheduling of the KL penalty coefficient $\beta$ was applied to control the KL divergence of the posterior distribution. Automatically, checkpoints of VAE models of different strengths were saved. For CMG, the training data used for VAE was 8 policies generated by CoMeDi on CMG. Due to a smaller amount of data, we used data augmentation to train the VAE.
Cooperator Agent The Cooperator agent was trained using PPO [29]. To promote exploration, the first 100M steps use reward shaping for dish and soup pick-up. We adapted our codebase on the previous implementation of HSP [41] and GAMMA [18].
GOAT For the GOAT policy training, we use the REINFORCE objective with KL divergence penalty for drifting away from the prior, mean, and standard deviation hard bounds to keep the policy constrained in the VAE latent space. This objective requires two separate passes of the environment to gather returns for self-play and cross-play.
Appendix C Hyperparameters
@@
The architecture of the policy model and the generative model are comparable. The representations are transformed into a variational posterior and action reconstruction predictions using an encoder head and a decoder head.
@@ Appendix D Computational Resources
Appendix E Minimax and Regret comparison on overcooked
Figure 7: Performance comparison between minimax vs GOAT objective on Multi-Strategy Counter when evaluated against held-out human behavior cloned model as partner. Error bars are the Standard Error of the Mean.
Appendix F Cooperative strategies learned by VAE
Figure 8: Left: We collect 4 cooperative policies on CMG-M using CoMeDi and then aggregate them. Right: The policies are used for training the VAE, where the latent dimension is 2 and we estimate the policy reconstruction from the VAE.
Figure 9: After training a VAE on a 4-population pool of CoMeDi agents, we visualize the generation by interpolating the dimensions of the latent vector. Particularly for this example, we choose the latent dimension to be 2. We can see that VAE trained on cooperative policies has only encoded those policies in the latent space. As we move from one side of the plot to the other, we can see that different cells are activated, and it covers the cooperative policies as seen in the Figure 8
Appendix G Bonferroni Tests
CoMeDi $1.93\times 10^{-11}$ $9.50\times 10^{-181}$
FCP $1.71\times 10^{-47}$ $2.24\times 10^{-53}$
BC $+$ RL $5.19\times 10^{-15}$ $1.20\times 10^{-20}$
MEP $1.0$ $3.48\times 10^{-28}$
MEP $+$ GAMMA $0.349$ $1.31\times 10^{-16}$
[…]
Appendix H Tests on CMG-S
(a) CMG-S layout
(b) Self-play
(c) TrajeDi
(d) MEP
(e) CoMeDi
(f) GOAT
Figure 10: We provide additional tests of the CMG-S problem space. In the problem space, rewards flow in a direction that creates a smooth learning landscape for the algorithms to exploit. The policies for the methods were aggregated and normalized into a heatmap. GOAT successfully assigns priorities to regions.
Appendix I Broader Impact Statement
Appendix J Human qualitative ratings.
(a) Counter circuit layout
(b) Multi-strategy Counter
Figure 11: Human evaluations of agent behavior. Participants rated their agreement with the statement "The agent adapted to me when making decisions" using a 5-point Likert scale ranging from Strongly Disagree to Strongly Agree. GOAT shows decent adaptability on simpler layout, whereas it appears more adaptable on harder layout.
(a) Counter circuit layout
(b) Multi-strategy Counter
Figure 12: User assessment of agent coordination. Participants rated their agreement with the statement "The agent’s actions were human-like" using a 5-point Likert scale ranging from Strongly Disagree to Strongly Agree. GOAT is seen as more confident in coordinating on harder layouts.
(a) Counter circuit layout
(b) Multi-strategy Counter
Figure 13: User assessments of agent frustration levels. Participants rated their agreement with the statement "The agent’s behavior was frustrating" using a 5-point Likert scale ranging from Strongly Disagree to Strongly Agree. GOAT receives comparable ratings for showing less frustrating behavior compared to their respective baselines.
Appendix K Limitation
The effectiveness of this method heavily depends on the training of VAE models and enough data available for VAE to learn the distribution instead of overfitting to a few examples. It would be a promising method for fields like robotics and human-AI interaction, where data availability is not an issue. We also found that in sparse reward problems, this method discovers solutions and trains the cooperator robustly.